# Past, Present, and Future of Citation Practices in HCI

### v1

@@ Abstract.
Science is a complex system comprised of many scientists who individually make collective decisions that, due to the size and nature of the academic system, largely do not affect the system as a whole. However, certain decisions at the meso-level of research communities, such as the Human-Computer Interaction (HCI) community, may result in deep and long-lasting behavioral changes in scientists. In this article, we provide evidence on how a change in editorial policies introduced at the ACM CHI Conference in 2016 launched the CHI community on an expansive path, denoted by a year-by-year increase in the mean number of references included in CHI articles. If this near-linear trend continues undisrupted, an article in CHI 2030 will include on average almost 130 references. Our meta-research provides insights into how the nature and meaning of citation practices in HCI have changed, influenced by factors such as digital accessibility of resources and academic pressures. The observed trend towards more citations reflects a citation culture where quantity is prioritized over quality, contributing to both author and peer reviewer fatigue. This article underscores the value of meta-research for research communities and the profound impact that meso-level policy adjustments have on the evolution of scientific fields and disciplines, urging stakeholders to carefully consider the broader implications of such changes.
journalyear: 2024
@@
[…]
Number of references per article in the Proceedings of the ACM CHI Conference from 1981 to 2024.
@@ 1. Introduction
Citation serves as the backbone of academic rigor, enabling researchers to build upon previous work, acknowledge contributions, and weave a rich tapestry of interdisciplinary knowledge. Editorial policies shape the landscape of academic publishing, influencing not only the structure and content of the research published but also the citation practices of researchers. Citation practices, as a tangible expression of scholarly discourse, provide valuable insights into the evolving norms and values of the CHI community. Understanding the impact of policy decisions on citation practices is crucial for ensuring that scholarly communication serves both the authors and the broader field effectively.
@@
In this work, we examine how the citation practices of researchers in the Human-Computer Interaction (HCI) community have changed after a significant editorial policy decision introduced in 2016. In this year, the page restrictions on references included in articles at the HCI field’s top conference, the ACM Conference on Human Factors in Computing Systems (CHI), were lifted. This policy decision had a profound impact on the citation practices of the CHI community. Our analysis of 11,542 articles published at the CHI Conference between 1981 and 2024 provides evidence of a year-by-year increase in the mean number of references included in CHI articles. This change signifies a departure from previous patterns and suggests the community’s entry into a novel trajectory (cf. Figure 1). If this trend continues undisrupted, an article in CHI 2030 will include on average about 130 references. Such a scenario poses practical challenges for authors and peer reviewers alike, straining the authors’ capacity to thoroughly engage with the literature and the peer reviewers’ capacity to vet the cited literature during the peer review process. It also raises fundamental questions about the nature of academic discourse and the balance between comprehensiveness and focus in scholarly communication.
@@
The excessive and escalating number of references in CHI articles raises important questions about the sustainability of current citation practices. The challenges brought by the citation practices within the HCI community are multifaceted, encompassing the expansion of the conference itself and the elevation of standards for what constitutes a publication-worthy contribution. Such developments, while indicative of progress, also contribute to the increasing complexity of the academic ecosystem, negatively impacting both authors and peer reviewers. With the volume of references expanding, the feasibility of conducting thorough and effective peer reviews is called into question, suggesting a pivotal moment for reflection and reevaluation of how the HCI field approaches the foundational practice of citation. In light of these findings, this paper aims to spark a conversation within the CHI community and beyond about the future of scholarly communication and academic publishing. By examining the contextual factors and ramifications of changing citation practices, we invite scholars, practitioners, and policymakers to reflect on past decisions and identify sustainable pathways forward.
2. Related Work
@@ 2.1. The Field of Human-Computer Interaction
Human-Computer Interaction is both a field and a discipline (Liu et al., 2014; Blackwell, 2015). As a field, HCI encompasses diverse disciplines, such as computer science, cognitive science, psychology, design, sociology, anthropology, and more (Kim, 1995; Blackwell, 2015; Liu et al., 2014; Oppenlaender and Hämäläinen, 2023). HCI draws from these diverse areas to understand how humans interact with computers and how to design user-friendly systems. As an academic discipline, HCI is taught in structured curricula at many universities in dedicated degree programs, educating students in the principles and practices of HCI. Human-Computer Interaction is also recognized as a professional discipline with specific roles, such as UX researchers and interaction designers. Other researchers have aimed to look past the dichotomy of field versus discipline by framing HCI as problem-solving (Oulasvirta and Hornbæk, 2016). In the remainder of this paper, we refer to HCI as a field.
@@
Each year, a growing number of articles are being published in the field of HCI. The international top conference in the field of HCI is the ACM CHI Conference on Human Factors in Computing Systems (CHI), which has been held each year since 1981 (with exception of 1984). CHI is the most prestigious conference venue for researchers in HCI, with many HCI labs submitting their articles only to this conference venue. As such, the CHI conference can serve as a reference standard for the entire field and discipline of HCI field. From its humble beginnings in the early 1980s, the conference has been expanding since 2005 in terms of articles accepted at the conference (see Figure 2). Except for the COVID-19 pandemic years 2021 and 2022, the CHI proceedings have exhibited near linear growth since 2005. In 2024, the CHI Proceedings exceeded 1000 articles. While this expansion is plan-driven and directed by CHI’s Steering Committee, it also reflects the expansion of the HCI field.
Figure 2. Number of research articles published in the CHI proceedings in the years 1981 to 2024 (excluding session details, panel sessions, and abstract-only entries).
[…]
Number of CHI articles from 1981 to 2024.
@@
In this expansive and evolving research landscape, the practice of citing relevant works is important for HCI researchers. Citations serve as a vital link, connecting current research efforts with the rich history of past works. Citations acknowledge the contributions of other researchers and weave a rich tapestry of knowledge that informs and enriches new inquiries. For a field as interdisciplinary as HCI, where the integration of diverse perspectives and methodologies is crucial, citations provide a structured way to navigate and summarize the vast array of existing knowledge for readers without deep subject-matter knowledge and expertise. Further, citations enhance the credibility and rigor of academic discourse, enabling scholars to build upon a verified body of work. Proper citation practices contribute to the integrity of the HCI field, fostering an environment of collaboration and continuous learning. As such, citations are not merely a scholarly obligation but a critical element that underpins the academic and intellectual growth of the HCI community.
@@ 2.2. Meta-Research on the CHI Community and the Field of HCI
While the majority of HCI researchers conduct research within their field, some researchers have also examined aspects about the field of HCI. This is commonly referred to as meta-research (Ioannidis et al., 2015). Meta-research is valuable, as it enables the research community to reflect on its research practices (Pochat, 2024). The CHI proceedings, in particular, are a suitable and fruitful study subject for meta-scientific and bibliometric inquiries in the field of HCI. Bartneck and Hu, for instance, presented a scientometric analysis of the CHI proceedings, focusing on organizations and countries that contribute to CHI (Bartneck and Hu, 2009). Their work highlighted the difficulty of judging quality in the context of best paper awards, finding a mismatch between awarded papers and citations received. Our paper also includes an analysis of awards, however we focus on analyzing whether there is a systematic bias in awarded papers.
Lee et al. presented a bibliometric analysis of the CHI proceedings (1981 to 2018, excluding 1984), using a citation network analysis to identify emerging topics in CHI (Lee et al., 2019). Liu et al. also conducted a bibliometric analysis of CHI publications. Their keyword co-word analysis quantified and described the thematic evolution of the HCI field based on 3152 CHI articles published between 1994 and 2013. Like our work, their work includes a comparison between two time periods (1994–2003, 2004–2013), highlighting the underlying trends in the HCI community.
Pohl and Mottelson presented a quantitative meta-research study on paper writing at CHI, including an investigation on citations received. Their analysis of 6578 CHI papers found that readability, title length, and novelty influence citation counts. Our work differs in that we do not investigate writing, and only focus on citations given, not received. Another area of meta-research is the use of literature reviews in HCI. Stefanidi et al. conducted an analysis of CHI proceedings and ACM Transactions on Computer—Human Interaction (TOCHI) articles included 189 articles (Stefanidi et al., 2023). Like in our work, the authors note an “insufficient consensus of what to expect of literature reviews in HCI.” The diverse types and methodologies of literature reviews in HCI prompt us to split our analysis of literature reviews into a broad and narrow part.
Kaye presented a quantitative analysis of CHI, examining author counts, gender, and repeat authorship (Kaye, 2009). Among other findings, the work identified a trend toward an increasing number of authors per paper. We identify the same trend in our work, and similar to Kaye, we aim to encourage a discussion about about what the preferred state of CHI should be.
@@ 3. Method
3.1. Dataset Collection
@@
The ACM CHI Conference on Human Factors in Computing Systems is the HCI field’s largest and most important annual conference. Therefore, the CHI conference proceedings are a fertile ground for investigating the citation practices of HCI authors. The CHI proceedings are stored in ACM’s Digital Library (ACM–DL), which is the main data source in our research.
@@
We collected data on the number of references for each article published in the CHI proceedings from 1981 to 2024111With exception of the year 1984 in which the CHI conference was not scheduled. by scraping the ACM–DL. We excluded session details, panel sessions, keynotes, and abstract-only entries using regular expressions. For articles with more than 150 references, we manually collected the number of references due to the ACM–DL’s display limit. Additionally, we identified and corrected six instances where the number of references was incorrectly reported as zero. This process resulted in a dataset of DOI identifiers with their respective reference counts. We then downloaded the PDF documents for each DOI from the ACM–DL, resulting in a comprehensive dataset of 11,542 CHI articles.

### v7

@@ Abstract.
Science is a complex system comprised of many scientists who individually make decisions that, due to the size and nature of the academic system, largely do not affect the system as a whole. However, certain decisions at the meso-level of research communities, such as the Human-Computer Interaction (HCI) community, may result in deep and long-lasting behavioral changes in scientists. In this article, we provide empirical evidence on how a change in editorial policies introduced at the ACM CHI Conference in 2016 destabilized the CHI research community and launched it on an expansive path, denoted by a year-by-year increase in the mean number of references included in CHI articles. If this near-linear trend continues undisrupted, an article at CHI 2030 will include on average almost 130 references. The trend toward more citations reflects a citation culture where quantity is prioritized over quality, contributing to both author and peer reviewer fatigue. Our exploratory analysis highlights the profound impact of meso-level policy adjustments on the evolution of scientific fields and disciplines, urging all stakeholders to carefully consider the broader implications of such changes.
references, citations, CHI, bibliometric analysis, event study, meta-science, meta-HCI
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713556isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Human computer interaction (HCI)ccs: General and reference Surveys and overviews
@@ 1. Introduction
Citation serves as the backbone of academic rigor, enabling researchers to build upon previous work, acknowledge contributions, and weave a rich tapestry of interdisciplinary knowledge. Citation practices, as a tangible expression of scholarly discourse, provide valuable insights into the evolving norms and values of the research community (Oppenlaender and Hosio, 2025). Editorial policies shape the landscape of academic publishing, influencing not only the structure and content of the research published but also the citation practices of researchers. Understanding the impact of policy decisions on citation practices is crucial for ensuring that scholarly communication serves both the authors and the broader field effectively.
In this work, we examine how the citation practices of researchers in the Human-Computer Interaction (HCI) community have changed after an editorial policy decision introduced in 2016. In this year, the page restrictions at HCI’s top conference, the ACM Conference on Human Factors in Computing Systems (CHI), were lifted. This policy decision had a profound impact on the citation practices of the CHI community. We examine the impact of this policy decision with three research questions. We first explore the historic development of the number of references in CHI articles:
  • RQ1:
    How has the number of citations from CHI articles to other works developed in the CHI Proceedings?
@@
Our analysis of 11,542 articles published at the ACM CHI Conference between 1981 and 2024 provides empirical evidence of a year-by-year increase in the mean number of references included in CHI articles. This trend was enabled by the change in policies in 2016, allowing authors to include an unlimited number of references in their articles. The subsequent change in the community’s citation practices signifies a departure from previous patterns and suggests the community’s entry into a novel trajectory (cf. Figure 1). If this trend continues undisrupted, articles at CHI 2030 will include on average about 130 references. Such a scenario poses practical challenges for authors and peer reviewers alike, straining the authors’ capacity to meaningfully engage with the literature and the peer reviewers’ capacity to thoroughly vet the cited literature during the peer review process.
We explore a number of potential reasons for the observed increase in the mean number of references included in CHI articles:
  • RQ2:
    What other factors, beside the 2016 policy decision, could potentially contribute to the observed growth in references, and how have these factors evolved over the years?
We plot and and visually analyze a number of co-occurring trends. As part of this, we also investigate whether there is a bias in awarding articles at CHI post-policy change. If articles with a high number of references were to systematically be presented with awards at the CHI Conference, this could send a signal to the community and potentially incentivize authors to include more references in their articles, contributing to the observed growth in the mean number of references per CHI article.
Last, we investigate the significance of these observed trends in shaping the expansive citation practices in the CHI community:
  • RQ3:
    Was there a significant change in citation practices of authors at CHI after the editorial policy change in 2016? If so, how do the co-occuring factors contribute to this trend?
We investigate this research question with an event study, demonstrating that the policy change at CHI ’16 had a profound and destabilizing effect on the CHI community. The CHI community has clearly entered a different trajectory at this point in time (cf. Figure 1), marked by increasing academic pressures contributing to fatigue of both authors and peer reviewers.
@@
The rising number of references in CHI articles raises important questions about the nature of academic discourse, the balance between comprehensiveness and focus in scholarly communication, and the sustainability of current citation practices. The challenges brought by the escalating citation practices within the HCI community are multifaceted, encompassing the expansion of the CHI Conference, the diversification of topics and methodologies, and the elevation of standards for what constitutes a publication-worthy contribution. Such developments, while indicative of progress, also contribute to the increasing complexity of the academic ecosystem, negatively impacting both authors and peer reviewers. With the volume of references expanding, the feasibility of conducting thorough and effective peer reviews is called into question. This article discusses a number of potential solutions to this trend.
We argue it is high time for the CHI community to pause, reflect, and reevaluate how the HCI field approaches the practice of citation. In light of our findings, this paper aims to spark a conversation within the CHI community and beyond about the future of scholarly communication and academic publishing. By examining the contextual factors and ramifications of changing citation practices, we invite scholars, practitioners, and policymakers to reflect on past decisions and identify sustainable pathways forward.
2. Background and Related Work
@@ 2.1. The Field of Human-Computer Interaction
Human-Computer Interaction is both a field and a discipline (Liu et al., 2014; Blackwell, 2015). As a field, HCI encompasses diverse disciplines, such as computer science, cognitive science, psychology, design, sociology, anthropology, and more (Kim, 1995; Blackwell, 2015; Liu et al., 2014; Oppenlaender and Hämäläinen, 2023). HCI draws from these diverse areas to understand how humans interact with computers and how to design user-friendly systems. As an academic discipline, HCI is taught in structured curricula at many universities in dedicated degree programs, educating students in the principles and practices of HCI. Human-Computer Interaction is also recognized as a professional discipline with specific roles, such as UX researchers and interaction designers. In the remainder of this paper, we refer to HCI as a field.
@@
Each year, a growing number of articles are being published in the field of HCI. The international top conference in HCI is the ACM Conference on Human Factors in Computing Systems (CHI), which has been held each year since 1981 (with exception of 1984). CHI is the most prestigious conference venue for researchers in HCI, with many labs submitting their articles exclusively to only this conference venue. As such, the CHI Conference can serve as a reference standard for the entire field and discipline of HCI.
From its humble beginnings in the early 1980s, the CHI Conference has been expanding since 2005 in terms of articles accepted at the conference. Except for the COVID-19 pandemic years in 2021 and 2022, the CHI Proceedings have exhibited near linear growth since 2005. In 2024, the CHI Proceedings exceeded 1,000 articles. While this expansion is plan-driven and directed by CHI’s Steering Committee, it also reflects the expansion and diversification of the wider field of HCI.
@@
In this expansive and evolving research landscape, the practice of citing relevant works is important for HCI researchers. Citations serve as a vital link, connecting current research efforts with the rich history of past works. Citations acknowledge the contributions of other researchers and weave a rich tapestry of knowledge that informs and enriches new inquiries. For a field as interdisciplinary as HCI, where the integration of diverse perspectives and methodologies is crucial, citations provide a structured way to navigate and summarize the vast array of existing knowledge for readers without deep subject-matter knowledge and expertise. Further, citations enhance the credibility and rigor of academic discourse, enabling scholars to build upon a verified body of work. Proper citation practices contribute to the integrity of the HCI field, fostering an environment of collaboration and continuous learning. As such, citations are not merely a scholarly obligation but a critical element in the academic and intellectual growth of the HCI community.
@@ 2.2. Meta-Research on the CHI Community and the Field of HCI
While the majority of HCI researchers conduct research within their field, some researchers have also examined aspects about the field of HCI. This is commonly referred to as meta-research (Ioannidis et al., 2015). Meta-research is valuable, as it enables the research community to reflect on its research practices (Pochat, 2024; Oppenlaender et al., 2025). The CHI Proceedings, in particular, are a fruitful study subject for meta-scientific and bibliometric inquiries in the field of HCI, due to the size and importance of the CHI Conference for the field of HCI. Bartneck and Hu, for instance, presented a scientometric analysis of the CHI Proceedings, focusing on organizations and countries that contribute to CHI (Bartneck and Hu, 2009). Their work highlighted the difficulty of judging quality in the context of best paper awards, finding a mismatch between awarded papers and citations received. Our paper also includes an analysis of awards, however we focus on analyzing whether there is a systematic bias in awarded papers.
Lee et al. presented a bibliometric analysis of the CHI Proceedings (from 1981 to 2018), using a citation network analysis to identify emerging topics in CHI (Lee et al., 2019). Liu et al. also conducted a bibliometric analysis of CHI publications (Liu et al., 2014). Their keyword co-word analysis quantified and described the thematic evolution of the HCI field based on 3,152 CHI articles published between 1994 and 2013. Like our work, their article includes a comparison between two time periods (1994–2003, 2004–2013), highlighting the underlying trends in the HCI community.
Pohl and Mottelson presented a quantitative meta-research study on paper writing at CHI, including an investigation on citations received (Pohl and Mottelson, 2019). Their analysis of 6,578 CHI papers found that readability, title length, and novelty influence citation counts. Our work differs in that we do not investigate writing and only focus on citations given, not received.
Another area of meta-research is the use of literature reviews in HCI. Stefanidi et al. conducted an analysis of the CHI Proceedings and ACM Transactions on Computer-Human Interaction (TOCHI) including 189 articles (Stefanidi et al., 2023). Like in our work, the authors note an “insufficient consensus of what to expect of literature reviews in HCI.” The diverse types and methodologies of literature reviews in HCI prompt us to split our analysis of literature reviews into a broad and narrow part.
Finally, Kaye presented a quantitative analysis of CHI, examining author counts, gender, and repeat authorship (Kaye, 2009). Among other findings, the work identified a trend toward an increasing number of authors per paper. We identify the same trend in our work, and similar to Kaye, we aim to encourage a discussion about what the preferred state of CHI should be.
@@ 3. Method
In this paper, we employ both exploratory analysis (with descriptive statistics and visual analysis) and statistical analysis with linear regression models and an event study specification. The research involved gathering, quantifying, and analyzing data from the ACM Digital Library (ACM–DL), as detailed in the following section.
3.1. Data Collection
@@
The ACM Conference on Human Factors in Computing Systems (CHI) is the HCI field’s largest and most important annual conference. Therefore, the Proceedings of the CHI Conference are a proxy for investigations into the wider field of HCI. In particular, the digital proceedings are a fertile ground for investigating the citation practices of HCI authors. The CHI Proceedings are stored in ACM’s Digital Library, which is the main source of data in our research.
@@
We collected all articles published in the CHI Proceedings from 1981 to 2024 (with exception of 1984) from the ACM–DL. We focused on full research articles and excluded articles in the companion proceedings or extended abstracts. We further excluded session details, panel sessions, keynotes, and abstract-only entries using regular expressions. For each article, we collected the full set of references included in the article and the set of authors from the ACM–DL. Further, we downloaded the PDF document and collected the full reference section from the ACM–DL for each article. This resulted in a large dataset of CHI articles ($N={11,542}$), with their respective full set of references and authors.
3.2. Analyzing the Mean Number of References in CHI Articles (RQ1)
To answer RQ1, we count the number of references in each CHI article. For each year in the CHI Proceedings, we calculate the mean, median, minimum, and maximum number of references. Visual inspection yields the insight that a clear change in slope occurred after the year 2015 (cf. Figure 1).
Simple linear regression analysis was conducted to evaluate the extent to which Year could predict the mean and minimum number of References after the policy change (2016–2024). We report the regression equation for each model, R2 values explaining the variance in the models, t-values and p-values for the slope coefficient, and the error metrics MAE (mean absolute error) and MSE (mean squared error). The t-values for slope quantify how many standard errors the estimated slope coefficient ($\beta_{1}$) is away from zero. It is used to test the null hypothesis that the slope is zero, which would imply no linear relationship between the predictor (Year) and the response variable (References). F-tests are conducted testing the overall significance of the regression models, reporting F-statistic and p-value. The two linear regression models are then used to predict the mean and minimum number of citations in future CHI proceedings.
The mean number of references per CHI article is the key variable in our work. In the following section, we describe how we visualized and analyzed a number of other contextual factors potentially playing a role in shaping the citation practices of the CHI community.
3.3. Exploring Other Contextual Factors (RQ2)
To answer RQ2, we explore the historical development of a number of co-occurring trends.
3.3.1. Analyzing bias in awarding articles
We examine whether a systematic bias can be observed in awarded papers in past proceedings of the CHI Conference. As award, we consider both “best paper” awards and “honorable mentions.” We count the references in awarded and unawarded articles at the CHI Conference. We then statically analyze the difference in the mean number of references in articles that received an award (Mawarded) and the mean number of references in articles that did not receive an award (Mregular) for each proceedings year from 2007 to 2024. We report the t-test statistic, p-values, 95% confidence intervals, level of significance, and effect size (using Cohen’s $d$).
3.3.2. Collaboration and co-authorship at CHI
For each article in the dataset, we count the number of authors. We plot the development of co-authorship over time and visually analyze the correlation of growth trends in authors and publications at CHI.
3.3.3. Literature reviews
Literature reviews aim to capture the current state of a research area or field. Given the rising number of scholarly articles (at CHI and globally), we expect that CHI authors are increasingly conducting literature surveys. To study this, we identified literature reviews by searching each CHI article’s full text (excluding the references section) for the keywords “literature review,” “literature survey,” and “systematic literature review.” We manually verified each occurrence, ensuring that only articles explicitly stating having conducted a literature review were included. Note that CHI authors, particularly in earlier proceedings, often referred to the related work section as a literature review. These instances are included in our plots. Additionally, some authors claim to have conducted an “extensive” or “exhaustive” literature review without detailing the process. Our data encompasses “literature review” in a broad sense, while “systematic literature reviews” are represented more narrowly, typically indicating articles that adhere to a more rigorous scientific process. We visualize the results, testing our assumption of a rise in the absolute and relative number of literature reviews at CHI.
3.3.4. Citations to pre-print articles
The landscape of academic publishing is changing. We examine how CHI authors cite pre-print articles. Pre-print articles are versions of scholarly articles that precede publication in peer-reviewed journals and conference venues. While traditionally shunned upon due to not having undergone formal peer review, it has become common place in some research fields, such as Machine Learning, to cite pre-print articles. One of the most popular pre-print archives is arXiv, launched by Cornell University in 1991. In recent years, an exponentially growing number of articles are being submitted to this pre-print archive. The high number of available pre-prints and the success of pre-prints in other fields may motivate CHI authors to also include more citations to pre-prints in their CHI articles.
To investigate this, we identified the number of citations to arXiv pre-prints in the references of each article in the CHI Proceedings by counting the number of references including a mention of “arxiv.” For counting the number of occurrences of arXiv in the references section, we used a case-insensitive regular expression that accounts for hyphenation within the word arXiv due to potential line breaks. We then plot the absolute and relative number of citations to arXiv pre-prints for each proceedings year and visually analyze the results.
3.3.5. Citations to datasets and software code repositories
With the rise of data-driven research methodologies and machine learning, CHI papers may have started referencing more data sources, scientific software and tools, code repositories, and prior data-heavy studies, which could be another potential factor contributing to the increase in the mean number of references in CHI articles. We investigate this trend by counting the number of references to five popular open data and source code repositories:
  • osf.io, a cloud-based management solution for open science, often used to share datasets and other data related to HCI studies,
As above, we limit our investigation to the articles’ reference section. We ignore citations in footnotes within the article’s full text to make the results comparable to the other findings in this article. We plot the number of citations to data and source code repositories and visually analyze the results.
3.3.6. Citations to questionable publishers
Predatory journals are questionable publications that accept articles with minimal peer review. These journals are generally regarded as having a lower quality compared to more rigorously peer-reviewed journals (Beall, 2012). Without trade-off cost between references in CHI articles, CHI authors may cite articles from potential predatory journals and questionable publishers more often.
To investigate this, we compare the references included in CHI articles with entries in Beall’s List of Potential Predatory Journals and Publishers (Beall, 2012). We first downloaded Beall’s List [^1] https://beallslist.net and merged the lists of journals and publishers into one list. We removed any additional notes and entries that are generic, such as “Qualitative Research” and “American Journal,” since these entries would produce many false positive matches. We then checked for exact matches between the entries in Beall’s List and the references included in CHI articles (1981–2024), using basic string normalization techniques (e.g., lower-casing, removal of line breaks and special characters, and expansion of ampersand characters into the word ’and’). This comparison yielded 5,126 potential matches between entries in Beall’s List and CHI references. Because the names of predatory journals are often designed to partially match with reputable journals (Beall, 2012), each of the 5,126 potential matches was manually reviewed to sort out false positives. We plot the number of citations to potential predatory journals and publishers and visually analyze the results.
Note that this approach is limited in several ways, pertaining to the limitations of Beall’s List and the string matching approach. First, Beall’s List is subjective and it should not be the only source of quality appraisal. Second, the list contains entries with notes that the journal or publisher may, in fact, not be predatory. Beall’s List also mentions several instances where a previously reputable journal has been taken over by a predatory publisher. Third, while an archived version is available, Beall’s List is no longer maintained. Finally, exact string matching is imperfect, although we address false positives with manual review. The presented figures can be considered a lower boundary, as there may be more matches (especially if including generic titles) that are not identified with the exact matching approach.
3.4. Analyzing the Significance of the Observed Trend (RQ3)
To answer RQ3, we analyze the impact of the policy change on the outcome variable (mean number of references in CHI articles) with an orthogonalized event study specification. An event study is a statistical method used to assess the impact of an event or intervention on an outcome by analyzing changes in the outcome variable before and after the event within a defined time window. The event of interest, here, is the 2016 policy decision to lift page restrictions on the references included in CHI articles, and the event windows are the equal-sized pre- and post-event periods (2007–2015 and 2016–2024). With the event study, we examine both the significance of the observed change in citation behavior among CHI authors and the effect of different confounders on the mean number of references in CHI articles. The rich set of covariates includes for each year $t$:
  • $\textrm{Authors}_{t}$: mean number of authors per article,
  • $\textrm{Arxiv}_{t}$: mean number of citations to arXiv pre-prints per article,
  • $\textrm{Repos}_{t}$: mean number of citations to code and data repositories per article,
  • $\textrm{Reviews}_{t}$: mean number of literature reviews per article,
  • $\textrm{Predators}_{t}$: mean number of citations to potential predatory journals and questionable publishers per article, and
  • $\textrm{AwardRatio}_{t}$: the relative prominence of awarded articles compared to unawarded articles, defined as the ratio between the mean number of references ($M$) in awarded and unawarded articles in a given proceedings year $t$:
    $\textrm{AwardRatio}_{t}=M_{\textrm{awarded},t}/M_{\textrm{unawarded},t}$.
The regression model for the event study can be written as:
$\textrm{References}_{t}=\beta_{0}+\beta_{1}\textrm{Post}_{t}+\beta_{2}\textrm{AwardRatio}_{t}+\beta_{3}\textrm{Authors}_{t}~{}+$
$\beta_{4}\textrm{Arxiv}_{t}+\beta_{5}\textrm{Repos}_{t}+\beta_{6}\textrm{Reviews}_{t}+\beta_{7}\textrm{Predators}_{t}~{}+$
$\beta_{8}(\textrm{Post}_{t}\cdot t)+\beta_{t}t+\epsilon_{t}$
with $\epsilon_{t}$ being an unobserved error term. $\textrm{Post}_{t}$ is an indicator variable denoting the time period (1 if $t\geq 2016$ and 0 otherwise). The parameters $\beta_{1}$ and $\beta_{8}$ are the key coefficients of interest in this event study specification. Coefficient $\beta_{1}$ indicates a mean difference in References between the pre- and post-event periods. The interaction term $\beta_{8}$ captures a change in slope at the event point ($t=2015.5$). The term $\beta_{t}$ captures trends that vary linearly over time, ensuring that the model accounts for systematic changes that occur regardless of the intervention.
The time variable $t$ is centered around its mean (the event point, $t=2015.5$). Orthogonalization was applied to address multicollinearity among predictors, ensuring that the estimated coefficients reflect independent contributions of each variable. That is, we regressed each variable on $t$ and use only the residuals $u_{t}$ for regression:
$\textrm{Variable}_{t}=\gamma_{0}+\gamma_{1}t+u_{t}$
where $u_{t}$ is the part of the variable not explained by $t$. This approach retains variability in the levels of variables and removes the temporal correlation while keeping other meaningful relationships. We report the coefficients, t-statistic for each coefficient with p-values, 95% confidence intervals, and overall significance of the model with an F-test.